OpenStudy (anonymous):

Score: Which set of side lengths will form a right triangle? 6, 8, 11 10, 24, 25 7, 24, 26 8, 15, 17

Directrix (directrix):

6, 8, 11 -------- Take the longest side 11 Square it --> 121 Get the sum of the squares of 6 and 8 6^2 + 8^2 = 36 + 64 = 100 Is 100 = 121 NO, so 6, 8, 11 will not form the sides of a right triangle. @Griffin16 Use the same technique (converse of Pythagorean Theorem) to determine if any of the remaining triples can be sides of a right triangle.